x1..Use the distributive property to remove the parentheses.
9(v – 2)

2.
Holly paid 7% sales tax and a disposal charge of $16 when she bought new tires for her car. The total cost was $259.96. If x represents the price of the new set of tires, then x satisfies x + 0.07x + 16 = 259.96. Find the price of her new set of tires.

3.
Solve for T: PV = nRT
Thank you!

Answer by rfer(12657)

1)
9(v-2)
multiply through by 9
9v-18
2)
x+0.07x+16=259.96
1.07x+16=259.96
subtract 16 from both sides.
1.07x=259.95-16
1.07x=243.96
divide both sides by 1.07
x=243.96/1.07
x=228 cost of tires
3)
PV=nRT
Divide both sides by nR
T=PV/nR